Please help!
If each side has a scale factor of 1" to 2', the actual perimeter is ____ and the actual area is _____.
Thank You!

Answer by venugopalramana(3286)

Please help!
If each side has a scale factor of 1" to 2',
1"=2*12"=24"...SO RATIO IS 24"/1"=24.HENCE
the actual perimeter is 24 TIMES THE SCALE LENGTH____ and the actual area is _24*24=576 TIMES THE SCALE AREA.____.
Thank You!